# Large Language Model Unlearning via Embedding-Corrupted Prompts

### v1

@@ D.1 Prompt Classifier Thresholding
We see that prompt classifiers without thresholding already perform well on most datasets. An exception is the out-of-distribution WMDP, where we have few samples for questions from one threat category. Specifically, it has a high false negative rate for forget samples and a non-trivial false positive rate on retain samples. Increasing the threshold in simple-threshold classifiers reduces the false positive rate to near-perfect, but the false negative rate also increases. By employing conformal prediction, we successfully reduce the false negative rate by more than 50%. In practice, we recommend selecting the thresholding method based on its performance on a relatively large held-out set to balance missing forget or retain samples. Depending on the risk posed by the unlearning target, conformal prediction might be a better choice in high-risk scenarios to reduce false positive predictions.
@@ D.2 Corruption Function Variants
In the experiments below, we do not tune the corruption strength for each corruption function but use the same corruption strength for similar functions. For example, for all random noise corruption, we use the same corrupted strength as the one picked in the main paper (Figure 2). In Table 9, we show that sign flipping and zero-out have consistently high forget quality, while the randomized corruption function might require extra strength tuning, especially for larger forget sets. This suggests that tuning the corruption strength based on the criteria used is important to achieve unlearning.
@@
In Tables 10 and 11, we present the results of eight variants of the corruptions on BBC News and HP Book unlearning. We see that most corruption functions used can achieve an ASG score below 5 while maintaining low perplexity and high unique token ratio.
@@
We see that the selection of corruption function and corruption strength is not as important on unlearning BBC News and HP Book as it is on the TOFU dataset, based on text similarity metrics. This suggests that the forget quality metric is a more rigorous measure than mere text similarity, as it evaluates the distributional similarity between the outputs of an unlearned model and a retrained model. Therefore, in practice, we recommend always searching for the best corruption function and corruption strength based on the available criteria.
@@ E.1 Time Delays of Prompt Content Detection
In Table 12, we report the per-example time delay (in milliseconds) introduced by running the prompt classifier during inference of the main LLM. The time in w/o Classifier(s), w/ Classifier(s), and Abs. Increase columns are measured on a machine with a single NVIDIA A100 using a Llama-2-7B(-Chat) with a batch size of 4. Note that the prompt content detection step only depends on the incoming prompt and is agnostic to the LLM subject to unlearning, so the statistics in the table is constant with respect to any other LLMs given fixed prompts. The largest delay of 85 ms is from TOFU (Forget10), which involves extra inference time by an additional token classifier based on BERT to identify tokens that are names in the prompt. In most cases, the extra delay is no more than 21 ms.
@@ E.2 TOFU
The results in this section are supporting evidence for Section 4.2 and Figure 2 in the main paper. We include the full results of Llama-2-7B-Chat and Phi-1.5 in Table 13 and Table 14, respectively. These results encompass all metrics described in Section C.1.1: conditional probability of the answer given the prompts, truth ratio (TR), ROUGE-L scores, model utility, and forget quality. We report all scores for the retain set, forget set, real authors, and world facts across all forget set sizes. In Figure 2, we plot the model utility and forget quality as shown in the last two columns of each table.
Besides the random noise and zero-out variants of ECO, we also include a sign-flip variant. This variant flips the signs of all entries in the embedding vectors of the selected tokens. In both tables, Table 13 and Table 14, the sign-flip variant exhibits low forget quality across all splits. This outcome likely stems from its higher (better) truth ratio compared to the retained model, leading to substantially different distributions from those of the truth ratio in the retained model. We hypothesize that this effect arises from the drastic alterations in the embedding vectors caused by flipping their signs.
@@
E.2.1 TOFU Example Generations Under Embedding Corruptions
@@ E.3 WMDP
In Figure 4, we visualize the average WMDP accuracy versus the model size. We see that effectiveness of unlearning using a prompting baseline decreases as the original performance of the model increases. For ECO, the accuracy after unlearning does not dependent on their original performance.
@@ E.3.1 Probing Evaluation
In Figure 3, we showcase the linear probe’s test accuracy in recovering the correct choice based on output logits from Zephyr-7B, Yi-34B-Chat, and Mixtral-8x7B-Instruct. Before using ECO to unlearn, a substantial proportion of the labels can to recovered by the linear probe classifier for all three models. After incorporating ECO in the forward pass, the classifier’s accuracy drops to random-chance level, indicating the effectiveness of ECO against recovering knowledge from the logit space.
@@ E.4 MMLU
The results presented in this section serve as supporting evidence for Section 4.3 and Table 2. We include the full results of 68 models on MMLU subsets unlearning in Table 18. For all models, ECO results in minimal to no performance loss on the corresponding retain subject, attributed to the prompt classifier’s low false positive rate.
Note that the forget accuracy for economics remains at 35.8 across multiple models. We manually inspected the predictions of these models and found that the corrupted prompts bias the predictions towards answer D. Given that the correct answers for the economics questions in MMLU are not uniformly distributed, with about 35% being D, the answers is still considered as random-guessing. Therefore, the universal effectiveness of our method is maintained.
In Figure 4, we visualize the average MMLU subset accuracy versus the model size. The pattern observed on MMLU subsets remains the same on that of the WMDP benchmark: while prompting could reduce the performance by a lot in some cases, the unlearned model maintains high accuracy. ECO consistently reduce the accuracy to random-guessing across all model sizes.
@@
Baichuan2-13B-Chat [134] 54.5 44.4 43.0 35.1 71.3 27.8 59.1 53.3 43.2 39.1 35.1 74.1 30.8 55.5 56.7 55.5 24.6 24.0 24.4 35.1 71.3 27.8 59.1
Baichuan2-7B-Chat [134] 48.4 40.3 36.5 25.4 65.7 29.2 55.3 48.6 40.5 35.7 25.4 64.8 29.5 53.5 54.5 54.0 29.2 24.6 24.2 25.4 65.7 29.2 55.3
c4ai-command-r-plus-4bit [28] 73.6 61.0 57.6 63.2 82.4 32.7 68.8 72.8 58.2 54.9 63.2 81.5 33.8 64.3 63.8 59.2 20.8 27.2 23.6 63.2 82.4 32.7 68.8
c4ai-command-r-v01-4bit [28] 65.9 55.8 49.8 52.6 86.1 33.5 66.2 64.5 55.9 49.0 52.6 87.0 33.2 58.6 58.9 57.8 20.7 24.8 25.0 52.6 86.1 33.5 66.2
dbrx-instruct [29] 76.1 59.5 54.9 59.6 82.4 42.4 70.4 75.0 57.9 49.4 59.6 79.6 33.0 63.9 64.7 56.4 22.6 24.5 24.9 59.6 82.4 42.4 70.4
deepseek-llm-67b-chat [48] 73.4 56.0 50.8 53.5 83.3 34.1 65.7 71.2 54.5 52.0 53.5 81.5 35.9 62.6 63.7 61.1 21.7 28.2 23.3 53.5 83.3 34.1 65.7
deepseek-llm-7b-chat [48] 43.8 40.4 36.5 31.6 62.0 30.0 59.0 41.1 40.1 32.8 31.6 59.3 28.9 54.5 56.8 54.3 26.9 27.5 24.5 31.6 62.0 30.0 59.0
deepseek-moe-16b-chat [48] 43.9 38.4 33.0 33.3 61.1 24.1 59.1 43.8 38.6 30.7 33.3 60.2 24.9 54.2 55.2 54.4 35.8 23.7 28.2 33.3 61.1 24.1 59.1
DeepSeek-V2-Chat [30] 84.9 60.6 60.9 65.8 89.8 44.3 66.8 79.8 38.6 55.9 65.8 82.4 44.6 62.8 59.3 57.2 21.6 24.2 27.8 65.8 89.8 44.3 66.8
DeepSeek-V2-Lite-Chat [30] 61.1 43.0 41.0 40.4 64.8 34.6 62.1 59.7 37.9 41.6 41.2 69.4 33.2 55.4 55.8 54.4 33.0 25.3 28.4 40.4 64.8 34.6 62.1
falcon-180B-chat [4] 67.2 53.4 42.6 40.4 77.8 31.4 62.7 66.2 52.9 40.4 40.4 77.8 30.0 60.8 60.9 59.6 20.5 24.7 23.4 40.4 77.8 31.4 62.7
falcon-40b-instruct [4] 54.3 42.5 35.5 27.2 63.0 27.3 62.4 52.9 41.0 34.0 27.2 61.1 27.8 56.6 56.5 55.2 26.4 23.7 22.2 27.2 63.0 27.3 62.4
gemma-1.1-2b-it [121] 40.0 36.7 29.1 21.1 50.9 30.0 54.3 40.0 36.1 28.7 21.1 47.2 28.4 48.0 48.8 48.6 22.6 24.0 26.1 21.1 50.9 30.0 54.3
gemma-1.1-7b-it [121] 65.9 45.8 44.9 42.1 72.2 35.4 61.4 63.1 45.1 43.9 42.1 70.4 33.5 49.4 48.5 47.7 22.3 27.7 22.7 42.1 72.2 35.4 61.4
gemma-2b-it [121] 35.7 33.2 27.9 24.6 49.1 24.3 52.5 33.9 32.3 28.5 24.6 42.6 23.0 45.9 46.0 45.9 26.6 25.7 27.7 24.6 49.1 24.3 52.5
gemma-7b-it [121] 48.2 38.9 39.1 33.3 63.9 30.8 58.8 50.0 39.1 39.5 33.3 63.0 30.0 51.9 51.8 50.8 21.6 26.3 23.5 33.3 63.9 30.8 58.8
@@
internlm2-chat-20b [20] 50.6 34.0 40.2 39.5 49.1 35.7 63.5 46.5 32.1 37.1 39.5 47.2 33.0 53.0 55.2 52.6 23.5 25.0 24.3 39.5 49.1 35.7 63.5
internlm2-chat-7b [20] 53.8 37.9 37.3 38.6 72.2 29.2 62.9 31.4 28.6 24.6 38.6 41.7 25.4 53.4 55.9 50.9 23.7 24.5 26.5 38.6 72.2 29.2 62.9
jetmoe-8b-chat [114] 42.0 38.7 35.2 34.2 53.7 24.1 56.6 34.7 36.3 29.5 31.6 46.3 25.7 50.2 50.4 50.0 23.8 24.4 21.8 31.6 53.7 25.1 56.6
Llama-2-13b-chat-hf [124] 49.5 42.8 35.2 25.4 65.7 29.2 59.1 47.1 40.5 34.4 25.4 65.7 28.1 48.8 49.6 49.5 27.3 26.6 25.1 25.4 65.7 29.2 59.1
Llama-2-70b-chat-hf [124] 64.0 49.5 45.1 38.6 80.6 32.7 61.0 59.4 44.0 43.6 38.6 72.2 31.4 49.5 50.4 48.1 20.5 24.7 23.4 38.6 80.6 32.7 61.0
Llama-2-7b-chat-hf [124] 38.1 37.7 31.1 29.8 53.7 28.6 55.7 36.0 34.8 27.0 29.8 51.9 25.7 45.7 45.5 45.1 22.6 27.8 23.7 29.8 53.7 28.6 55.7
Llama3-ChatQA-1.5-70B [81] 84.2 61.9 65.4 61.4 85.2 44.1 66.3 85.0 61.0 64.8 61.4 85.2 41.1 61.9 61.7 60.3 35.8 23.9 28.1 61.4 85.2 44.1 66.3
@@
Meta-Llama-3-70B-Instruct [3] 85.4 64.1 70.7 73.7 86.1 50.3 67.4 83.4 61.1 65.4 73.7 84.3 47.3 56.3 55.7 53.9 35.8 23.8 28.1 73.7 86.1 50.3 67.4
Meta-Llama-3-8B-Instruct [3] 71.2 50.2 47.3 48.2 77.8 38.6 62.8 66.7 46.8 50.0 48.2 77.8 36.5 50.3 51.1 50.2 20.4 24.1 24.3 48.2 77.8 38.6 62.8
Mistral-7B-Instruct-v0.1 [60] 54.1 42.3 38.9 31.6 67.6 33.0 60.7 53.0 39.8 38.9 31.6 65.7 31.9 54.4 54.9 53.8 26.7 24.5 26.9 31.6 67.6 33.0 60.7
Mistral-7B-Instruct-v0.2 [60] 60.2 45.1 43.6 48.2 72.2 34.1 62.7 35.7 29.8 21.1 48.2 27.8 28.9 48.4 46.9 45.7 24.2 24.0 27.0 48.2 72.2 34.1 62.7
Mistral-7B-Instruct-v0.3 [60] 59.2 48.7 43.4 47.4 75.9 35.7 64.5 57.8 42.7 40.4 45.6 73.1 33.2 55.1 52.0 51.9 26.8 24.0 26.0 47.4 75.9 35.7 64.5
Mixtral-8x22B-Instruct-v0.1 [61] 80.9 60.0 61.9 64.0 84.3 41.9 67.0 77.2 59.0 58.6 64.0 80.6 43.0 57.8 60.7 53.7 28.0 24.2 27.5 64.0 84.3 41.9 67.0
Mixtral-8x7B-Instruct-v0.1 [61] 71.3 54.8 54.7 57.0 79.6 39.5 66.2 50.6 51.5 43.2 57.0 78.7 34.1 55.9 58.5 54.2 20.6 24.7 23.4 57.0 79.6 39.5 66.2
OLMo-7B-Instruct-hf [45] 47.8 39.8 35.7 28.9 48.1 27.0 56.6 49.4 37.8 34.6 28.9 46.3 26.8 51.6 51.0 49.8 21.3 26.8 26.6 28.9 48.1 27.0 56.6
openchat-3.5-0106-gemma [129] 64.5 51.8 49.8 53.5 74.1 35.7 67.9 63.4 49.6 49.4 53.5 75.9 33.5 62.2 62.4 61.3 35.8 23.8 28.1 53.5 74.1 35.7 67.9
openchat-3.5-0106 [129] 64.0 50.9 45.1 43.9 75.9 35.1 66.9 62.7 46.8 45.1 43.9 75.0 32.2 53.0 54.2 50.7 19.8 23.2 24.1 43.9 75.9 35.1 66.9
openchat-3.6-8b-20240522 [129] 69.7 50.0 50.4 48.2 78.7 37.8 67.1 67.5 48.7 50.6 47.4 80.6 36.5 58.6 59.8 57.3 20.2 24.1 24.3 48.2 78.7 37.8 67.1
Orca-2-13b [93] 58.1 46.3 41.0 33.3 76.9 32.7 61.7 58.0 45.8 40.0 33.3 77.8 31.4 52.5 51.8 52.6 29.5 24.3 24.9 33.3 76.9 32.7 61.7
Orca-2-7b [93] 49.7 42.5 39.1 25.4 63.9 32.2 58.4 48.2 41.3 34.0 25.4 63.0 29.2 50.7 53.3 51.8 23.6 27.9 22.3 25.4 63.9 32.2 58.4
phi-1_5 [71] 41.2 34.2 33.8 30.7 51.9 29.5 56.0 42.0 34.8 34.0 30.7 48.1 26.8 54.4 54.8 54.6 22.4 23.9 25.9 30.7 51.9 29.5 56.0
@@
Phi-3-medium-128k-instruct [1] 73.1 56.4 56.4 57.9 78.7 32.2 64.3 80.7 56.0 61.1 57.9 86.1 37.6 52.9 54.5 52.9 21.7 24.7 24.7 57.9 78.7 32.2 64.3
Phi-3-medium-4k-instruct [1] 84.6 61.8 67.6 61.4 86.1 41.4 66.1 58.0 40.4 56.4 49.1 73.1 38.9 50.7 51.7 50.1 21.9 25.4 26.0 61.4 86.1 41.4 66.1
Phi-3-mini-128k-instruct [1] 70.4 50.7 49.0 49.1 66.7 32.4 62.3 48.4 46.3 41.4 43.0 61.1 30.8 45.1 45.6 45.1 24.4 26.5 25.7 49.1 66.7 32.4 62.3
Phi-3-mini-4k-instruct [1] 76.9 53.1 54.3 52.6 80.6 37.8 62.2 36.6 25.6 32.0 27.2 49.1 30.8 44.8 44.1 43.8 19.7 26.3 23.4 52.6 80.6 37.8 62.2
Phi-3-small-128k-instruct [1] 70.7 57.1 52.5 52.6 78.7 34.1 66.6 75.3 56.6 55.3 50.0 79.6 31.6 52.7 52.8 52.3 35.9 23.3 28.0 52.6 78.7 34.1 66.6
Phi-3-small-8k-instruct [1] 85.2 60.8 63.7 59.6 84.3 39.5 69.4 46.8 32.0 27.5 22.8 38.9 28.9 46.7 45.7 47.4 35.8 24.4 28.0 59.6 84.3 39.5 69.4
Qwen1.5-0.5B-Chat [8] 36.8 34.1 30.7 23.7 37.0 23.0 43.3 23.4 25.6 22.5 23.7 25.9 25.1 38.4 38.5 38.3 20.5 24.5 23.5 23.7 37.0 23.0 43.3

### v2

@@ D.1 Prompt Classifier Thresholding
We observe that prompt classifiers without thresholding already perform well on most datasets. An exception is the out-of-distribution WMDP, where we have few samples for questions from one threat category. Specifically, it has a high false negative rate for forget samples and a non-trivial false positive rate on retain samples. Increasing the threshold in simple-threshold classifiers reduces the false positive rate to near-perfect, but the false negative rate also increases. By employing conformal prediction, we successfully reduce the false negative rate by more than 50%. In practice, we recommend selecting the thresholding method based on its performance on a relatively large held-out set to balance missing forget or retain samples. Depending on the risk posed by the unlearning target, conformal prediction might be a better choice in high-risk scenarios to reduce false positive predictions.
@@ D.2 Corruption Function Variants
In the experiments below, we do not tune the corruption strength for each corruption function but use the same corruption strength for similar functions. For example, for all random noise corruption, we use the same corruption strength as the one selected in the main paper (Figure 2). In Table 12, we show that sign flipping and zero-out have consistently high forget quality, while the randomized corruption function might require extra strength tuning, especially for larger forget sets. This suggests that tuning the corruption strength based on the criteria used is important to achieve effective unlearning.
@@
In Tables 13 and 14, we present the results of eight variants of the corruptions on BBC News and HP Book unlearning. We see that most corruption functions used can achieve an ASG score below 5 while maintaining low perplexity and high unique token ratio.
@@
We see that the selection of the corruption function and corruption strength is not as important for unlearning BBC News and HP Book as it is for the TOFU dataset, based on text similarity metrics. This suggests that the forget quality metric is a more rigorous measure than mere text similarity, as it evaluates the distributional similarity between the outputs of an unlearned model and a retrained model. Therefore, in practice, we recommend always searching for the best corruption function and corruption strength based on the available criteria.
D.3 Task-Agnostic Selection of the Surrogate Metric Value
In section 3.3, we select the surrogate metric value, $\hat{v}_{r}$, for each task individually. While this serves as a sound experimental setup, it raises concerns regarding its adaptability to real-world use cases, where a metric might not be directly available or easy to calculate.
To address this, we performed an additional experiment using a dataset- or task-agnostic selection of $\hat{v}_{r}$. Specifically, we use Llama-3-70B-Instruct to generate 100 synthetic responses in one of two forms: 1) stating “I do not know the answer to the question” (IDK), or 2) refusal (e.g., “I cannot answer the question”). These responses are independent of the datasets and tasks under consideration. We also use the LLM subjected to unlearning to generate its original response (before unlearning). We then apply the four text similarity metrics used in the copyrighted content unlearning task to measure the difference between 1) the original responses and 2) the synthetic IDK or refusal responses. The goal is to minimize this difference across all three tasks.
In essence, we aim to push the model’s output toward IDK or refusal responses by using a zeroth-order objective to minimize the textual similarity between the model output and the template responses, regardless of the specific task.
Task Task-dependent selection of $\hat{v}_{r}$ Task-agnostic selection of $\hat{v}_{r}$
TOFU (forget quality $\uparrow$) 0.9674 0.9188
WMDP Chemistry (accuracy $\downarrow$) 26.6 25.8
BBC News (ASG $\downarrow$) 11.8 11.8
Table 15: Comparison of task-dependent and task-agnostic selection of $\hat{v}_{r}$ for all tasks. Note that the score in row 3 does not change because we used the same way to select $\hat{v}_{r}$ for copyrighted content tasks.
In table 15, we show that task-agnostic selection still maintains the effectiveness of unlearning, suggesting a task-dependent select of the surrogate metric value is not necessary.
@@ E.1 Time Delays of Prompt Content Detection
In Table 16, we report the per-example time delay (in milliseconds) introduced by running the prompt classifier during the inference of the main LLM. The times in the w/o Classifier(s), w/ Classifier(s), and Abs. Increase columns are measured on a machine with a single NVIDIA A100 using a Llama-2-7B(-Chat) with a batch size of 4. Note that the prompt content detection step only depends on the incoming prompt and is agnostic to the LLM subject to unlearning, so the statistics in the table are constant with respect to any other LLMs given fixed prompts. The largest delay of 85 ms is from TOFU (Forget10), which involves extra inference time by an additional token classifier based on BERT to identify tokens that are names in the prompt. In most cases, the extra delay is no more than 21 ms.
@@ E.2 TOFU
E.2.1 Full Results
The results in this section provide supporting evidence for Section 4.2 and Figure 2 in the main paper. We include the full results of Llama-2-7B-Chat and Phi-1.5 in Table 17 and Table 18, respectively. These results encompass all metrics described in Section C.1.1: conditional probability of the answer given the prompts, truth ratio (TR), ROUGE-L scores, model utility, and forget quality. We report all scores for the retain set, forget set, real authors, and world facts across all forget set sizes. In Figure 2, we plot the model utility and forget quality as shown in the last two columns of each table.
Besides the random noise and zero-out variants of ECO, we also include a sign-flip variant. This variant flips the signs of all entries in the embedding vectors of the selected tokens. In both Table 17 and Table 18, the sign-flip variant exhibits low forget quality across all splits. This outcome likely stems from its higher (better) truth ratio compared to the retained model, leading to substantially different distributions from those of the truth ratio in the retained model. We hypothesize that this effect arises from the drastic alterations in the embedding vectors caused by flipping their signs.
@@ E.2.1 Full Results
E.2.2 TOFU Example Generations Under Embedding Corruptions
@@ E.3.1 Full Results
In Figure 4, we visualize the average WMDP accuracy versus the model size. We observe that the effectiveness of unlearning using a prompting baseline decreases as the original performance of the model increases. For ECO, the accuracy after unlearning does not depend on the original performance.
@@ E.3.2 Probing Evaluation
In Figure 3, we showcase the linear probe’s test accuracy in recovering the correct choice based on output logits from Zephyr-7B, Yi-34B-Chat, and Mixtral-8x7B-Instruct. Before using ECO to unlearn, a substantial proportion of the labels can be recovered by the linear probe classifier for all three models. After incorporating ECO in the forward pass, the classifier’s accuracy drops to random-chance level, indicating the effectiveness of ECO in preventing knowledge recovery from the logit space.
@@ E.4 MMLU
E.4.1 Full Results
The results presented in this section serve as supporting evidence for Section 4.3 and Table 2. We include the full results of 68 models on MMLU subset unlearning in Table 22. For all models, ECO results in minimal to no performance loss on the corresponding retain subject, attributed to the prompt classifier’s low false positive rate.
Note that the forget accuracy for economics remains at 35.8 across multiple models. We manually inspected the predictions of these models and found that the corrupted prompts bias the predictions toward answer D. Given that the correct answers for the economics questions in MMLU are not uniformly distributed, with about 35% being D, the answers are still considered as random-guessing. Therefore, the universal effectiveness of our method is maintained.
In Figure 4, we visualize the average MMLU subset accuracy versus the model size. The pattern observed on MMLU subsets mirrors that of the WMDP benchmark: while prompting could significantly reduce performance in some cases, the unlearned model maintains high accuracy. ECO consistently reduces the accuracy to random-guessing across all model sizes.
@@ E.4.1 Full Results
Baichuan2-13B-Chat [141] 54.5 44.4 43.0 35.1 71.3 27.8 59.1 53.3 43.2 39.1 35.1 74.1 30.8 55.5 56.7 55.5 24.6 24.0 24.4 35.1 71.3 27.8 58.7
Baichuan2-7B-Chat [141] 48.4 40.3 36.5 25.4 65.7 29.2 55.3 48.6 40.5 35.7 25.4 64.8 29.5 53.5 54.5 54.0 29.2 24.6 24.2 25.4 65.7 29.2 55.2
c4ai-command-r-plus-4bit [29] 73.6 61.0 57.6 63.2 82.4 32.7 68.8 72.8 58.2 54.9 63.2 81.5 33.8 64.3 63.8 59.2 20.8 27.2 23.6 63.2 82.4 32.7 68.4
c4ai-command-r-v01-4bit [29] 65.9 55.8 49.8 52.6 86.1 33.5 66.2 64.5 55.9 49.0 52.6 87.0 33.2 58.6 58.9 57.8 20.7 24.8 25.0 52.6 86.1 33.5 66.1
dbrx-instruct [30] 76.1 59.5 54.9 59.6 82.4 42.4 70.4 75.0 57.9 49.4 59.6 79.6 33.0 63.9 64.7 56.4 22.6 24.5 24.9 59.6 82.4 42.4 70.1
deepseek-llm-67b-chat [51] 73.4 56.0 50.8 53.5 83.3 34.1 65.7 71.2 54.5 52.0 53.5 81.5 35.9 62.6 63.7 61.1 21.7 28.2 23.3 53.5 83.3 34.1 64.4
deepseek-llm-7b-chat [51] 43.8 40.4 36.5 31.6 62.0 30.0 59.0 41.1 40.1 32.8 31.6 59.3 28.9 54.5 56.8 54.3 26.9 27.5 24.5 31.6 62.0 30.0 58.2
deepseek-moe-16b-chat [51] 43.9 38.4 33.0 33.3 61.1 24.1 59.1 43.8 38.6 30.7 33.3 60.2 24.9 54.2 55.2 54.4 35.8 23.7 28.2 33.3 61.1 24.1 57.9
DeepSeek-V2-Chat [31] 84.9 60.6 60.9 65.8 89.8 44.3 66.8 79.8 38.6 55.9 65.8 82.4 44.6 62.8 59.3 57.2 21.6 24.2 27.8 65.8 89.8 44.3 65.1
DeepSeek-V2-Lite-Chat [31] 61.1 43.0 41.0 40.4 64.8 34.6 62.1 59.7 37.9 41.6 41.2 69.4 33.2 55.4 55.8 54.4 33.0 25.3 28.4 40.4 64.8 34.6 61.7
falcon-180B-chat [4] 67.2 53.4 42.6 40.4 77.8 31.4 62.7 66.2 52.9 40.4 40.4 77.8 30.0 60.8 60.9 59.6 20.5 24.7 23.4 40.4 77.8 31.4 62.5
falcon-40b-instruct [4] 54.3 42.5 35.5 27.2 63.0 27.3 62.4 52.9 41.0 34.0 27.2 61.1 27.8 56.6 56.5 55.2 26.4 23.7 22.2 27.2 63.0 27.3 61.7
gemma-1.1-2b-it [128] 40.0 36.7 29.1 21.1 50.9 30.0 54.3 40.0 36.1 28.7 21.1 47.2 28.4 48.0 48.8 48.6 22.6 24.0 26.1 21.1 50.9 30.0 53.9
gemma-1.1-7b-it [128] 65.9 45.8 44.9 42.1 72.2 35.4 61.4 63.1 45.1 43.9 42.1 70.4 33.5 49.4 48.5 47.7 22.3 27.7 22.7 42.1 72.2 35.4 61.2
gemma-2b-it [128] 35.7 33.2 27.9 24.6 49.1 24.3 52.5 33.9 32.3 28.5 24.6 42.6 23.0 45.9 46.0 45.9 26.6 25.7 27.7 24.6 49.1 24.3 52.2
gemma-7b-it [128] 48.2 38.9 39.1 33.3 63.9 30.8 58.8 50.0 39.1 39.5 33.3 63.0 30.0 51.9 51.8 50.8 21.6 26.3 23.5 33.3 63.9 30.8 58.4
@@
internlm2-chat-20b [21] 50.6 34.0 40.2 39.5 49.1 35.7 63.5 46.5 32.1 37.1 39.5 47.2 33.0 53.0 55.2 52.6 23.5 25.0 24.3 39.5 49.1 35.7 63.3
internlm2-chat-7b [21] 53.8 37.9 37.3 38.6 72.2 29.2 62.9 31.4 28.6 24.6 38.6 41.7 25.4 53.4 55.9 50.9 23.7 24.5 26.5 38.6 72.2 29.2 62.2
jetmoe-8b-chat [120] 42.0 38.7 35.2 34.2 53.7 24.1 56.6 34.7 36.3 29.5 31.6 46.3 25.7 50.2 50.4 50.0 23.8 24.4 21.8 31.6 53.7 25.1 56.2
Llama-2-13b-chat-hf [131] 49.5 42.8 35.2 25.4 65.7 29.2 59.1 47.1 40.5 34.4 25.4 65.7 28.1 48.8 49.6 49.5 27.3 26.6 25.1 25.4 65.7 29.2 58.6
Llama-2-70b-chat-hf [131] 64.0 49.5 45.1 38.6 80.6 32.7 61.0 59.4 44.0 43.6 38.6 72.2 31.4 49.5 50.4 48.1 20.5 24.7 23.4 38.6 80.6 32.7 60.9
Llama-2-7b-chat-hf [131] 38.1 37.7 31.1 29.8 53.7 28.6 55.7 36.0 34.8 27.0 29.8 51.9 25.7 45.7 45.5 45.1 22.6 27.8 23.7 29.8 53.7 28.6 55.5
Llama3-ChatQA-1.5-70B [84] 84.2 61.9 65.4 61.4 85.2 44.1 66.3 85.0 61.0 64.8 61.4 85.2 41.1 61.9 61.7 60.3 35.8 23.9 28.1 61.4 85.2 44.1 66.2
@@
Meta-Llama-3-70B-Instruct [3] 85.4 64.1 70.7 73.7 86.1 50.3 67.4 83.4 61.1 65.4 73.7 84.3 47.3 56.3 55.7 53.9 35.8 23.8 28.1 73.7 86.1 50.3 66.6
Meta-Llama-3-8B-Instruct [3] 71.2 50.2 47.3 48.2 77.8 38.6 62.8 66.7 46.8 50.0 48.2 77.8 36.5 50.3 51.1 50.2 20.4 24.1 24.3 48.2 77.8 38.6 62.5
Mistral-7B-Instruct-v0.1 [63] 54.1 42.3 38.9 31.6 67.6 33.0 60.7 53.0 39.8 38.9 31.6 65.7 31.9 54.4 54.9 53.8 26.7 24.5 26.9 31.6 67.6 33.0 59.7
Mistral-7B-Instruct-v0.2 [63] 60.2 45.1 43.6 48.2 72.2 34.1 62.7 35.7 29.8 21.1 48.2 27.8 28.9 48.4 46.9 45.7 24.2 24.0 27.0 48.2 72.2 34.1 62.5
Mistral-7B-Instruct-v0.3 [63] 59.2 48.7 43.4 47.4 75.9 35.7 64.5 57.8 42.7 40.4 45.6 73.1 33.2 55.1 52.0 51.9 26.8 24.0 26.0 47.4 75.9 35.7 64.3
Mixtral-8x22B-Instruct-v0.1 [64] 80.9 60.0 61.9 64.0 84.3 41.9 67.0 77.2 59.0 58.6 64.0 80.6 43.0 57.8 60.7 53.7 28.0 24.2 27.5 64.0 84.3 41.9 66.3
Mixtral-8x7B-Instruct-v0.1 [64] 71.3 54.8 54.7 57.0 79.6 39.5 66.2 50.6 51.5 43.2 57.0 78.7 34.1 55.9 58.5 54.2 20.6 24.7 23.4 57.0 79.6 39.5 64.8
OLMo-7B-Instruct-hf [48] 47.8 39.8 35.7 28.9 48.1 27.0 56.6 49.4 37.8 34.6 28.9 46.3 26.8 51.6 51.0 49.8 21.3 26.8 26.6 28.9 48.1 27.0 55.7
openchat-3.5-0106-gemma [136] 64.5 51.8 49.8 53.5 74.1 35.7 67.9 63.4 49.6 49.4 53.5 75.9 33.5 62.2 62.4 61.3 35.8 23.8 28.1 53.5 74.1 35.7 67.4
openchat-3.5-0106 [136] 64.0 50.9 45.1 43.9 75.9 35.1 66.9 62.7 46.8 45.1 43.9 75.0 32.2 53.0 54.2 50.7 19.8 23.2 24.1 43.9 75.9 35.1 65.5
openchat-3.6-8b-20240522 [136] 69.7 50.0 50.4 48.2 78.7 37.8 67.1 67.5 48.7 50.6 47.4 80.6 36.5 58.6 59.8 57.3 20.2 24.1 24.3 48.2 78.7 37.8 65.6
Orca-2-13b [98] 58.1 46.3 41.0 33.3 76.9 32.7 61.7 58.0 45.8 40.0 33.3 77.8 31.4 52.5 51.8 52.6 29.5 24.3 24.9 33.3 76.9 32.7 61.6
Orca-2-7b [98] 49.7 42.5 39.1 25.4 63.9 32.2 58.4 48.2 41.3 34.0 25.4 63.0 29.2 50.7 53.3 51.8 23.6 27.9 22.3 25.4 63.9 32.2 58.1
phi-1_5 [74] 41.2 34.2 33.8 30.7 51.9 29.5 56.0 42.0 34.8 34.0 30.7 48.1 26.8 54.4 54.8 54.6 22.4 23.9 25.9 30.7 51.9 29.5 55.7
@@
Phi-3-medium-128k-instruct [1] 73.1 56.4 56.4 57.9 78.7 32.2 64.3 80.7 56.0 61.1 57.9 86.1 37.6 52.9 54.5 52.9 21.7 24.7 24.7 57.9 78.7 32.2 63.9
Phi-3-medium-4k-instruct [1] 84.6 61.8 67.6 61.4 86.1 41.4 66.1 58.0 40.4 56.4 49.1 73.1 38.9 50.7 51.7 50.1 21.9 25.4 26.0 61.4 86.1 41.4 65.1
Phi-3-mini-128k-instruct [1] 70.4 50.7 49.0 49.1 66.7 32.4 62.3 48.4 46.3 41.4 43.0 61.1 30.8 45.1 45.6 45.1 24.4 26.5 25.7 49.1 66.7 32.4 62.1
Phi-3-mini-4k-instruct [1] 76.9 53.1 54.3 52.6 80.6 37.8 62.2 36.6 25.6 32.0 27.2 49.1 30.8 44.8 44.1 43.8 19.7 26.3 23.4 52.6 80.6 37.8 61.5
Phi-3-small-128k-instruct [1] 70.7 57.1 52.5 52.6 78.7 34.1 66.6 75.3 56.6 55.3 50.0 79.6 31.6 52.7 52.8 52.3 35.9 23.3 28.0 52.6 78.7 34.1 66.5
Phi-3-small-8k-instruct [1] 85.2 60.8 63.7 59.6 84.3 39.5 69.4 46.8 32.0 27.5 22.8 38.9 28.9 46.7 45.7 47.4 35.8 24.4 28.0 59.6 84.3 39.5 68.0
Qwen1.5-0.5B-Chat [8] 36.8 34.1 30.7 23.7 37.0 23.0 43.3 23.4 25.6 22.5 23.7 25.9 25.1 38.4 38.5 38.3 20.5 24.5 23.5 23.7 37.0 23.0 41.6